OpenStudy (anonymous):

Find the values of k for which the quadratic equation kx(x - 2) + 6 = 0 has two equal roots.

OpenStudy (anonymous):

for two equal roots, b^2 - 4ac = 0 simplifying we get kx^2 - 2kx + 6 = 0 so a = k b = -2k and c = 6 so (-2k)^2 - 4 (k)(6) = 0 4k^2 - 24k = 0 what to do next...??

OpenStudy (anonymous):

4k(k-6) = 0 k = 0 or k =6

OpenStudy (anonymous):

But is k = 0 a valid solution ???

OpenStudy (anonymous):

I think you might have become mixed up between the k's and the original equation in x. for k = 0 and for k = 6, the original expression in terms of x has two equal solutions.

OpenStudy (anonymous):

at least according to Coolsector... I think that's what he meant.

OpenStudy (anonymous):

k = 0 doesn't look valid to me though.

OpenStudy (anonymous):

kx(x - 2) + 6 = 0 for k = 6 (6)x(x-2) + 6 = 0 6x^2 - 12x + 6 = 0 6(x^2 - 2x + 1) = 0 (x-1)^2 = 0 two equal solutions are x = 1, x = 1

OpenStudy (anonymous):

@JakeV8 is right about it .. taking k = 0 wont do any good eventually after finding k =0 or k =6 we need to go back and see what they really mean k=0 mean 6=0 in the original equation so the only possible value is k=6

OpenStudy (anonymous):

yeah...that was my point...only 6 seems to be the valid solution here but the book said the values of k are 0 and 6 which is incorrect... Thanks guys for your time and effort....☺